OpenStudy (surryyy):

Two times the width of a certain rectangle exceeds its length by three inches, and eight times its length is twelve more than its perimeter. Find the width of the rectangle. Answer 6 inches 9 inches 12 inches 3 inches

OpenStudy (surryyy):

l = 2 w = 0

OpenStudy (midhun.madhu1987):

w cant be zero... it is the width of a rectangle...

OpenStudy (midhun.madhu1987):

Wait i made a mistake...

OpenStudy (midhun.madhu1987):

I got it.. wait

OpenStudy (midhun.madhu1987):

Let the Length be x and Breadth be y. Then , 2y = x + 3 x - 2y = 3 ------ (1) And 8x = 2(x + y) + 12 8x = 2x + 2y + 12 8x - 2x - 2y = 12 6x - 2y = 12 --- cancelling throughout by 2 we get 3x - y = 6 --------(2) Solve the 2 equations (1) & (2)

OpenStudy (midhun.madhu1987):

Did you understand it?

OpenStudy (nikato):

translate the two sentences. 2w=L +3 and 8L=12+2L+2w subtract both sides by 3 and you get L=2w-3 plug this into your second equation 8(2w-3)= 12+2(2w-3)+2w and you can solve for w, which is the width.

OpenStudy (surryyy):

w = 3

OpenStudy (surryyy):

the final answer is 3?

OpenStudy (nikato):

correct

OpenStudy (surryyy):

Thnx!